# A new HOC scheme for computation of flow and heat transfer in nonuniform polar grids

### v2

@@ Abstract
In this work, a higher order compact (HOC) discretization is developed on the nonuniform polar grid. The discretization conceptualized using the unsteady convection-diffusion equation (CDE) is further extended to flow problems governed by the Navies-Stokes (N-S) equations as well as the Boussinesq equations. The scheme developed here combines the advantages of body-fitted mesh with grid clustering, thereby making it efficient to capture flow gradients on polar grids. The scheme carries a spatial convergence of order three with temporal order of convergence being almost two. Diverse flow problems are being investigated using the scheme. Apart from two verification studies we validate the scheme by time marching simulation for the benchmark problem of driven polar cavity and the problem of natural convection in the horizontal concentric annulus. In the process, a one-sided approximation for the Neumann boundary condition for vorticity is also presented. Finally, the benchmark problem of forced convection around a circular cylinder is tackled. The results obtained in this study are analyzed and compared with the well-established numerical and experimental data wherever available in the literature. The newly developed scheme is found to generate accurate solutions in each case.
@@
Compact scheme , Navier-Stokes equation , Boussinesq equation , nonuniform polar grid , driven polar cavity , circular cylinder , natutral convection
@@ 1 Introduction
Fluid flow problems involving circular geometries have attracted a great deal of attention over the years owing to their theoretical significance and physical relevance. For such importance, higher order compact (HOC) schemes have been developed on the cylindrical polar coordinates. Preliminary works to develop HOC schemes on cylindrical polar coordinate systems mostly focused on the Poisson equation and uniform grids [1, 2, 3, 4, 5, 6]. Here it is important to note that in polar coordinates Poisson equation involves variable coefficients and thus the above-mentioned works involved significant endeavors and were not mere extension of the works done in Cartesian coordinates. Additionally, a valid discretization at singularity $r=0$ is necessary. Subsequently, considerable efforts could be found in the literature on HOC discretization of N-S equations in the body-fitted polar coordinate system [7, 8, 9, 10, 11]. In the year 2005 Sanyasiraju and Manjula [7] proposed a higher order semicompact scheme to solve the flow past an impulsively started circular cylinder. The authors used a wider stencil to discretize a few terms of the governing equation rendering semicompactness to the scheme. This strategy also helped alleviate challenges associated with the variable coefficients of the second order derivative. Subsequently, Kalita and Ray [8] developed a temporally second order and spatially third order accurate HOC scheme for streamfunction-vorticity $(\psi-\omega)$ formulation of unsteady N-S equation. As the authors worked with a modified differential equation the resulting difference equation carried a higher order of singularity. The scheme was specially adapted to simulate incompressible flow past a circular cylinder directly on nonuniform polar grids. A rather straightforward adaptation of this scheme for steady convection-diffusion equation (CDE) can be found in [9]. In the year 2013, the pioneering work on compact difference schemes for the pure streamfunction formulation of N-S equations in polar coordinates was reported by Yu and Tian [10]. The authors here worked with the steady biharmonic equation in polar coordinates. The scheme developed therein is of second order accuracy and carries streamfunction and its first derivatives as the unknown variables. Of late Das et al. [11] also worked with steady second order equation with variable coefficients in polar coordinates and introduced a third order accurate HOC scheme. The scheme claimed to be implemented on a nonuniform grid used an implicit form of first order derivatives and was successful in simulating steady incompressible flows. It is clear from the above discussion that classical HOC discretization of transient generalized second order PDE with variable coefficients in polar coordinates has not been attempted apart from the work of Kalita and Ray [8] on $\psi-\omega$ form of the N-S equations, whose extension to other governing equations such as the Boussinesq equation is not immediate. Additionally, it is intriguing to notice that developed HOC schemes in polar coordinates have been traditionally used to handle the nonuniformity of grids in radial direction only. However, heat and fluid flow problems are often associated with generation of steep gradients of streamfunction and vorticity near the bluff body or domain boundary walls, and as such grid clustering in all directions should be economical.
@@
In this work, we intend to develop a new transient HOC formulation that can simulate fluid as well as heat transfer problems directly in polar coordinates. The scheme conceived in the process must be generalizable to approximate second order equations with variable coefficients in polar coordinates and should be efficient for circular geometries. Starting with a second order PDE with variable coefficients, we advocate a discretization strategy amenable for extension to a nonlinear system of N-S equations and beyond. Additionally, we adopt the philosophy of nonuniform grids for accurate resolution of complex flow problems. We look forward to combining the virtues of Padé approximation of first order derivatives on nonuniform grids with variable coefficients and compact discretization of higher order derivatives in terms of flow variables and their first order gradients. Theoretically, the scheme has a temporal convergence of second order and a spatial convergence of order three. The scheme developed here leads to stable higher order discretization in conjunction with both Dirichlet and convective boundary conditions. The formulation is validated by applying it to the problem of an unsteady Gaussian pulse governed by the linear CDE. This problem is equipped with known analytical solution, which helps in error analysis. The spatial and temporal order of convergence are also established during this process. However, to comprehend the robustness and adaptability of this new scheme, we carry out the simulations for benchmark problems of both flow and heat transfer, namely the flow inside a driven polar cavity, natural convection inside a circular annulus, and forced convection around a stationary circular cylinder. These flows are governed by N-S equations and Boussinesq equations, both the equation being cast in transient incompressible form. The $(\psi-\omega)$ formulation of these coupled nonlinear PDE’s are conformed for the present computation.
@@ 2 Mathematical formulation and discretization procedure
$\partial_{r}\phi_{i,j}=\delta_{r}\phi_{i,j}-\frac{1}{2}\left(h_{r_{i}}-h_{r_{i-1}}\right)\partial_{rr}\phi_{i,j}-\frac{1}{6}\left(\frac{h_{r_{i}}^{3}+h_{r_{i-1}}^{3}}{h_{r_{i}}+h_{r_{i-1}}}\right)\partial_{rrr}\phi_{i,j}+\mathcal{O}\left(\frac{h_{r_{i}}^{4}-h_{r_{i-1}}^{4}}{h_{r_{i}}+h_{r_{i+1}}}\right),$ (4)
@@
$\partial_{rr}\phi_{i,j}=\delta_{rr}\phi_{i,j}-\frac{1}{3}\left(h_{r_{i}}-h_{r_{i-1}}\right)\partial_{rrr}\phi_{i,j}-\frac{1}{12}\left(\frac{h_{r_{i}}^{3}+h_{r_{i-1}}^{3}}{h_{r_{i}}+h_{r_{i-1}}}\right)\partial_{rrrr}\phi_{i,j}+\mathcal{O}\left(\frac{h_{r_{i}}^{4}-h_{r_{i-1}}^{4}}{h_{r_{i}}+h_{r_{i+1}}}\right).$ (5)
Combining equations 4 and 5, we arrive at the third order accurate approximation of second order derivative which is given as
$\partial_{rr}\phi_{i,j}=$ $2\left(\frac{2(h_{r_{i}}^{2}-h_{r_{i}}h_{r_{i-1}}+h_{r_{i-1}}^{2})}{h_{r_{i}}^{2}+h_{r_{i-1}}^{2}}\delta_{rr}\phi_{i,j}-\frac{h_{r_{i}}^{2}-h_{r_{i}}h_{r_{i-1}}+h_{r_{i-1}}^{2}}{h_{r_{i}}^{2}+h_{r_{i-1}}^{2}}\delta_{r}\phi_{r_{i,j}}\right.$ (6)
$\left.-\frac{h_{r_{i}}-h_{r_{i-1}}}{h_{r_{i}}^{2}+h_{r_{i-1}}^{2}}\delta_{r}\phi_{i,j}+\frac{h_{r_{i}}-h_{r_{i-1}}}{h_{r_{i}}^{2}+h_{r_{i-1}}^{2}}\phi_{r_{i,j}}\right)+\mathcal{O}\left(\frac{h_{r_{i}}^{4}-h_{r_{i-1}}^{4}}{h_{r_{i}}+h_{r_{i-1}}}\right).$
The analogous approximation along the tangential direction can be obtained as
$\partial_{\theta\theta}\phi_{i,j}=$ $2\left(\frac{2(h_{\theta_{j}}^{2}-h_{\theta_{j}}h_{\theta_{j-1}}+h_{\theta_{j-1}}^{2})}{h_{\theta_{j}}^{2}+h_{\theta_{j-1}}^{2}}\delta_{\theta\theta}\phi_{i,j}-\frac{h_{\theta_{j}}^{2}-h_{\theta_{j}}h_{\theta_{j-1}}+h_{\theta_{j-1}}^{2}}{h_{\theta_{j}}^{2}+h_{\theta_{j-1}}^{2}}\delta_{\theta}\phi_{\theta_{i,j}}\right.$ (7)
$\left.-\frac{h_{\theta_{j}}-h_{\theta_{j-1}}}{h_{\theta_{j}}^{2}+h_{\theta_{j-1}}^{2}}\delta_{\theta}\phi_{i,j}+\frac{h_{\theta_{j}}-h_{\theta_{j-1}}}{h_{\theta_{j}}^{2}+h_{\theta_{j-1}}^{2}}\phi_{\theta_{i,j}}\right)+\mathcal{O}\left(\frac{h_{\theta_{j}}^{4}-h_{\theta_{j-1}}^{4}}{h_{\theta_{j}}+h_{\theta_{j-1}}}\right).$
Once these approximations are achieved they are used in the CDE (1) to obtain its semi discrete form around the node $(i,j)$, which may be written as,
$a\left.\phi_{t}\right|_{i,j}+[\mathcal{A}\phi]_{i,j}=f_{i,j}.$ (8)
@@
$\mathcal{A}\phi_{i,j}=$ $-\left(2A_{1}\delta_{r}^{2}+\frac{2A_{2}}{r_{i}^{2}}\delta_{\theta}^{2}+\frac{A_{3}}{h_{r_{i-1}}}\delta_{r}+\frac{A_{4}}{r_{i}^{2}h_{\theta_{j-1}}}\delta_{\theta}\right)\phi_{i,j}$ (9)
$+\left(A_{1}\delta_{r}+\frac{A_{3}}{h_{r_{i-1}}}+c_{1}\right)\phi_{r_{i,j}}+\left(A_{2}\delta_{\theta}+\frac{A_{4}}{r_{i}^{2}h_{\theta_{j-1}}}+\frac{c_{2}}{r_{i}^{2}}\right)\phi_{\theta_{i,j}}$
which is accompanied by the coefficients,
$A_{1}=\frac{2(1-\alpha_{r}+\alpha_{r}^{2})}{(1+\alpha_{r}^{2})},\quad A_{2}=\frac{2(1-\alpha_{\theta}+\alpha_{\theta}^{2})}{(1+\alpha_{\theta}^{2})},\quad A_{3}=\frac{2(1-\alpha_{r})}{(1+\alpha_{r}^{2})},\quad A_{4}=\frac{2(1-\alpha_{\theta})}{(1+\alpha_{\theta}^{2})}.$
In the coefficients $A_{1}$, $A_{2}$, $A_{3}$ and $A_{4}$ the terms $\alpha_{r}$ and $\alpha_{\theta}$ are used to replace the arrays $\alpha_{r_{i}}$, $i\in{1,2,3,\dots,n_{r}-2}$ and $\alpha_{\theta_{j}}$, $j\in{1,2,3,\dots,n_{\theta}-2}$ respectively, which are characterized as follows,
$\alpha_{r_{i-1}}$ $=h_{r_{i}}/h_{r_{i-1}},\quad i\in\{2,3,\dots,n_{r}-1\},$
$\alpha_{\theta_{j-1}}$ $=h_{\theta_{j}}/h_{\theta_{j-1}},\quad j\in\{2,3,\dots,n_{\theta}-1\}.$
It is fairly evident that the values of $\alpha_{r}$ and $\alpha_{\theta}$ varies with the change in the values of grid spacing and so are the values of $A_{i}$’s, $i\in{1,2,3,4}$. Since we seek to time march from $n$-th to $(n+1)$-th level, we further carry out the temporal discretization of the unsteady term in equation 8. With the aid of Crank-Nicolson approximation, the fully discretized form of CDE (1) can be expressed as
$\left(1+\frac{\delta t}{2a}\mathcal{A}\right)\phi_{i,j}^{n+1}=\left(1-\frac{\delta t}{2a}\mathcal{A}\right)\phi_{i,j}^{n}+\frac{\delta t}{2a}\left(f_{i,j}^{n+1}+f_{i,j}^{n}\right).$ (10)
Finally, making use of the operators $\delta_{r}$, $\delta_{\theta}$, $\delta_{rr}$ and $\delta_{\theta\theta}$ in $\mathcal{A}$, the algebraic system associated with equation 10 can be derived as
@@
$-B_{6}\left(\phi_{r_{i+1,j}}^{n+1}-\phi_{r_{i-1,j}}^{n+1}\right)-\left(B_{7}+c_{1}\right)\phi_{r_{i,j}}^{n+1}$
$-B_{8}\left(\phi_{\theta_{i,j+1}}^{n+1}-\phi_{\theta_{i,j-1}}^{n+1}\right)-\left(B_{9}+\frac{c_{2}}{r_{i}^{2}}\right)\phi_{\theta_{i,j}}^{n+1}$
$-B_{6}\left(\phi_{r_{i+1,j}}^{n}-\phi_{r_{i-1,j}}^{n}\right)-\left(B_{7}+c_{1}\right)\phi_{r_{i,j}}^{n}$
$-B_{8}\left(\phi_{\theta_{i,j+1}}^{n}-\phi_{\theta_{i,j-1}}^{n}\right)-\left(B_{9}+\frac{c_{2}}{r_{i}^{2}}\right)\phi_{\theta_{i,j}}^{n}$

### v3

# A new transient higher order compact scheme for computation of flow and heat transfer in nonuniform polar grids

@@ Abstract
In this work, a higher order compact (HOC) discretization is developed on the nonuniform polar grid. The discretization conceptualized using the unsteady convection-diffusion equation (CDE) is further extended to flow problems governed by the Navies-Stokes (N-S) equations as well as the Boussinesq equations. The scheme developed here combines the advantages of body-fitted mesh with grid clustering, thereby making it efficient to capture flow gradients on polar grids. The scheme carries a spatial convergence of order three with temporal order of convergence being almost two. Diverse flow problems are being investigated using the scheme. Apart from a verification studies we validate the scheme by time marching simulation for the benchmark problem of driven polar cavity and the problem of natural convection in the horizontal concentric annulus. In the process, a one-sided approximation for the Neumann boundary condition for vorticity is also presented. Finally, the benchmark problem of forced convection around a circular cylinder is tackled. The results obtained in this study are analyzed and compared with the well-established numerical and experimental data wherever available in the literature. The newly developed scheme is found to generate accurate solutions in each case.
@@
Compact scheme , Navier-Stokes equation , Boussinesq equation , nonuniform polar grid , driven polar cavity , circular cylinder , natural convection
@@ 1 Introduction
Fluid flow problems involving circular geometries have attracted a great deal of attention over the years owing to their theoretical significance and physical relevance. For such importance, HOC schemes have been developed on the cylindrical polar coordinates. Preliminary works to develop HOC schemes on cylindrical polar coordinate systems mostly focused on the Poisson equation and uniform grids [1, 2, 3, 4, 5, 6]. Here it is important to note that in polar coordinates Poisson equation involves variable coefficients and thus the above-mentioned works involved significant endeavors and were not mere extension of the works done in the Cartesian coordinates. Additionally, a valid discretization at singularity $r=0$ is necessary. Subsequently, considerable efforts could be found in the literature on HOC discretization of N-S equations in the body-fitted polar coordinate system [7, 8, 9, 10, 11]. In the year 2005 Sanyasiraju and Manjula [7] proposed a higher order semicompact scheme to solve the flow past an impulsively started circular cylinder. The authors used a wider stencil to discretize a few terms of the governing equation rendering semicompactness to the scheme. This strategy also helped alleviate challenges associated with the variable coefficients of the second order derivative. Subsequently, Kalita and Ray [8] developed a temporally second order and spatially third order accurate HOC scheme for streamfunction-vorticity $(\psi-\omega)$ formulation of unsteady N-S equations. As the authors worked with a modified differential equation the resulting difference equation carried a higher order of singularity. The scheme was specially adapted to simulate incompressible flow past a circular cylinder directly on nonuniform polar grids. A rather straightforward adaptation of this scheme for steady CDE can be found in [9]. In the year 2013, the pioneering work on compact difference schemes for the pure streamfunction formulation of N-S equations in polar coordinates was reported by Yu and Tian [10]. The authors here worked with the steady biharmonic equation in polar coordinates. The scheme developed therein is of second order accuracy and carries streamfunction and its first derivatives as the unknown variables. Of late Das et al. [11] also worked with steady second order equation with variable coefficients in polar coordinates and introduced a third order accurate HOC scheme. The scheme claimed to be implemented on a nonuniform grid used an implicit form of first order derivatives and was successful in simulating steady incompressible flows. It is clear from the above discussion that classical HOC discretization of transient generalized second order PDE with variable coefficients in polar coordinates has not been attempted apart from the work of Kalita and Ray [8] on $\psi-\omega$ form of the N-S equations, whose extension to other governing equations such as the Boussinesq equation is not immediate. Additionally, it is intriguing to notice that developed HOC schemes in polar coordinates have been traditionally used to handle the nonuniformity of grids in radial direction only. However, heat and fluid flow problems are often associated with generation of steep gradients of streamfunction and vorticity near the bluff body or domain boundary walls, and as such grid clustering in all directions should be economical.
@@
In this work, we intend to develop a new transient HOC formulation that can simulate fluid as well as heat transfer problems directly in polar coordinates. The scheme conceived in the process must be generalizable to approximate second order equations with variable coefficients in polar coordinates and should be efficient for circular geometries. Starting with a second order PDE with variable coefficients, we advocate a discretization strategy amenable for extension to a nonlinear system of N-S equations and beyond. Additionally, we adopt the philosophy of nonuniform grids for accurate resolution of complex flow problems. We look forward to combining the virtues of Padé approximation of first order derivatives on nonuniform grids with variable coefficients and compact discretization of higher order derivatives in terms of flow variables and their first order gradients. Theoretically, the scheme has a temporal convergence of second order and a spatial convergence of order three. The scheme developed here leads to stable higher order discretization in conjunction with both Dirichlet and convective boundary conditions. The formulation is validated by applying it to the problem of an unsteady Gaussian pulse governed by the linear CDE. This problem is equipped with known analytical solution, which helps in error analysis. The spatial and temporal order of convergences are also established during this process. However, to comprehend the robustness and adaptability of this new scheme, we carry out the simulations for benchmark problems of both flow and heat transfer, namely the flow inside a driven polar cavity, natural convection inside a circular annulus, and forced convection around a stationary circular cylinder. These flows are governed by N-S equations and Boussinesq equations, both the equations being cast in transient incompressible form. The $\psi-\omega$ formulation of these coupled nonlinear PDE’s are conformed for the present computation.
@@ 2 Mathematical formulation and discretization procedure
$\partial_{r}\phi_{i,j}=\delta_{r}\phi_{i,j}-\frac{1}{2}\left(h_{r_{i}}-h_{r_{i-1}}\right)\partial_{rr}\phi_{i,j}-\frac{1}{6}\left(\frac{h_{r_{i}}^{3}+h_{r_{i-1}}^{3}}{h_{r_{i}}+h_{r_{i-1}}}\right)\partial_{rrr}\phi_{i,j}$ $-\frac{1}{24}\left(\frac{h_{r_{i}}^{4}-h_{r_{i-1}}^{4}}{h_{r_{i}}+h_{r_{i-1}}}\right)\partial_{rrrr}\phi_{i,j}$ (4)
$+\mathcal{O}\left(\frac{h_{r_{i}}^{5}+h_{r_{i-1}}^{5}}{h_{r_{i}}+h_{r_{i-1}}}\right),$
@@
$\partial_{rr}\phi_{i,j}=\delta_{r}^{2}\phi_{i,j}-\frac{1}{3}\left(h_{r_{i}}-h_{r_{i-1}}\right)\partial_{rrr}\phi_{i,j}-\frac{1}{12}\left(\frac{h_{r_{i}}^{3}+h_{r_{i-1}}^{3}}{h_{r_{i}}+h_{r_{i-1}}}\right)\partial_{rrrr}\phi_{i,j}$ $-\frac{1}{60}\left(\frac{h_{r_{i}}^{4}-h_{r_{i-1}}^{4}}{h_{r_{i}}+h_{r_{i-1}}}\right)\partial_{rrrrr}\phi_{i,j}$ (5)
$+\mathcal{O}\left(\frac{h_{r_{i}}^{5}+h_{r_{i-1}}^{5}}{h_{r_{i}}+h_{r_{i-1}}}\right).$
Substituting $\partial_{rrr}\phi_{i,j}$ from equation 4 in equation 5, we can obtain the following second order approximation for the second-order derivative given by
$\partial_{rr}\phi_{i,j}=\frac{h_{r_{i}}^{2}-h_{r_{i}}h_{r_{i-1}}+h_{r_{i-1}}^{2}}{h_{r_{i}}h_{r_{i-1}}}\delta_{r}^{2}\phi_{i,j}-\frac{2(h_{r_{i}}-h_{r_{i-1}})}{h_{r_{i}}h_{r_{i-1}}}(\delta_{r}\phi_{i,j}-\phi_{r_{i,j}})$ $-\frac{1}{12}h_{r_{i}}h_{r_{i-1}}\partial_{rrrr}\phi_{i,j}$ (6)
$+\mathcal{O}\left((h_{r_{i-1}}-h_{r_{i}})h_{r_{i}}h_{r_{i-1}}\right).$
It is easy to see that equation 6 reverts back to standard central approximation of order two on uniform grid.
Further, equation 4 yeilds
$\partial_{rr}\phi_{i,j}=\delta_{r}\phi_{r_{i,j}}-\frac{1}{2}\left(h_{r_{i}}-h_{r_{i-1}}\right)\partial_{rrr}\phi_{i,j}-\frac{1}{6}\left(\frac{h_{r_{i}}^{3}+h_{r_{i-1}}^{3}}{h_{r_{i}}+h_{r_{i-1}}}\right)\partial_{rrrr}\phi_{i,j}$ $-\frac{1}{24}\left(\frac{h_{r_{i}}^{4}-h_{r_{i-1}}^{4}}{h_{r_{i}}+h_{r_{i-1}}}\right)\partial_{rrrrr}\phi_{i,j}$ (7)
$+\mathcal{O}\left(\frac{h_{r_{i}}^{5}+h_{r_{i-1}}^{5}}{h_{r_{i}}+h_{r_{i-1}}}\right).$
Equations 5 and 7, when simplified give rise to the difference operator $(2\delta_{r}^{2}\phi_{i,j}-\delta_{r}\phi_{r_{i,j}})$ that carries a first-order truncation error for the second-order derivative and is given by
$\partial_{rr}\phi_{i,j}=(2\delta_{r}^{2}\phi_{i,j}-\delta_{r}\phi_{r_{i,j}})-\frac{1}{6}(h_{r_{i}}-h_{r_{i-1}})\partial_{rrr}\phi_{i,j}+\mathcal{O}\left(\frac{h_{r_{i}}^{4}-h_{r_{i-1}}^{4}}{h_{r_{i}}+h_{r_{i-1}}}\right).$ (8)
This necessitates a further approximation of $\partial_{rrr}\phi_{i,j}$ in equation 8 and is carried out using equation 4 to arrive at
$\partial_{rr}\phi_{i,j}=$ $2\left(\frac{h_{r_{i}}^{2}-h_{r_{i}}h_{r_{i-1}}+h_{r_{i-1}}^{2}}{h_{r_{i}}^{2}+h_{r_{i-1}}^{2}}(2\delta_{r}^{2}\phi_{i,j}-\delta_{r}\phi_{r_{i,j}})\right.$ (9)
$\left.-\frac{h_{r_{i}}-h_{r_{i-1}}}{h_{r_{i}}^{2}+h_{r_{i-1}}^{2}}(\delta_{r}\phi_{i,j}-\phi_{r_{i,j}})\right)+\frac{1}{12}(h_{r_{i}}-h_{r_{i-1}})^{2}\partial_{rrrr}\phi_{i,j}$
$+\mathcal{O}\left(\frac{(h_{r_{i}}-h_{r_{i-1}})(2h_{r_{i}}^{4}-2h_{r_{i}}^{3}h_{r_{i-1}}+3h_{r_{i}}^{2}h_{r_{i-1}}^{2}-2h_{r_{i}}h_{r_{i-1}}^{3}+2h_{r_{i-1}}^{4})}{(h_{r_{i}}^{2}+h_{r_{i-1}}^{2})}\right).$
Here, we note that the above second-order discretization not only attains optimal fourth-order convergence on uniform grid but also its leading truncation error term $\frac{1}{12}(h_{r_{i}}-h_{r_{i-1}})^{2}\partial_{rrrr}\phi_{i,j}$, as seen in equation 7, converges to zero quadratically as we move from nonuniform to uniform grid. Next, eliminating the fourth-order derivative from equations 6 and 9, we arrive at the following discretization of the second-order derivative with third-order truncation term on nonuniform grid as,
$\partial_{rr}\phi_{i,j}=$ $2\left\{\left(\frac{2h_{r_{i}}h_{r_{i-1}}}{h_{r_{i}}^{2}+h_{r_{i-1}}^{2}}+\frac{(h_{r_{i}}-h_{r_{i-1}})^{2}}{2h_{r_{i}}h_{r_{i-1}}}\right)\delta_{r}^{2}\phi_{i,j}-\frac{h_{r_{i}}h_{r_{i-1}}}{h_{r_{i}}^{2}+h_{r_{i-1}}^{2}}\delta_{r}\phi_{r_{i,j}}\right.$ (10)
$\left.-\frac{(h_{r_{i}}-h_{r_{i-1}})}{h_{r_{i}}^{2}-h_{r_{i}}h_{r_{i-1}}+h_{r_{i-1}}^{2}}\left(\frac{h_{r_{i}}h_{r_{i-1}}}{h_{r_{i}}^{2}+h_{r_{i-1}}^{2}}+\frac{(h_{r_{i}}-h_{r_{i-1}})^{2}}{h_{r_{i}}h_{r_{i-1}}}\right)(\delta_{r}\phi_{i,j}-\phi_{r_{i,j}})\right\}$
$+\mathcal{O}\left(\frac{h_{r_{i}}h_{r_{i-1}}(h_{r-i}^{3}-h_{r_{i}}^{3})}{(h_{r_{i}}^{2}+h_{r_{i-1}}^{2})}\right).$
It is quite evident that on uniform grid the above approximation shall revert back to compact fourth-order approximation. Compared to discretization in equation 9, approximation in equation 10 posses a higher-order truncation term and the order of approximation is optimal on nonuniform compact stencil.
Again following a similar course of action in the $\theta$-direction we get
$\partial_{\theta\theta}\phi_{i,j}=$ $2\left\{\left(\frac{2h_{\theta_{j}}h_{\theta_{j-1}}}{h_{\theta_{j}}^{2}+h_{\theta_{j-1}}^{2}}+\frac{(h_{\theta_{j}}-h_{\theta_{j-1}})^{2}}{2h_{\theta_{j}}h_{\theta_{j-1}}}\right)\delta_{\theta}^{2}\phi_{i,j}-\frac{h_{\theta_{j}}h_{\theta_{j-1}}}{h_{\theta_{j}}^{2}+h_{\theta_{j-1}}^{2}}\delta_{\theta}\phi_{\theta_{i,j}}\right.$ (11)
$\left.-\frac{(h_{\theta_{j}}-h_{\theta_{j-1}})}{h_{\theta_{j}}^{2}-h_{\theta_{j}}h_{\theta_{j-1}}+h_{\theta_{j-1}}^{2}}\left(\frac{h_{\theta_{j}}h_{\theta_{j-1}}}{h_{\theta_{j}}^{2}+h_{\theta_{j-1}}^{2}}+\frac{(h_{\theta_{j}}-h_{\theta_{j-1}})^{2}}{h_{\theta_{j}}h_{\theta_{j-1}}}\right)(\delta_{\theta}\phi_{i,j}-\phi_{\theta_{i,j}})\right\}$
$+\mathcal{O}\left(\frac{h_{\theta_{j}}h_{\theta_{j-1}}(h_{\theta_{j}}^{3}-h_{\theta_{j-1}}^{3})}{(h_{\theta_{j}}^{2}+h_{\theta_{j-1}}^{2})}\right).$
Making use of these spatial approximations, the CDE (1) around the node $(i,j)$ can be expressed in the following semidiscrete form,
$a\left.\frac{\partial\phi}{\partial t}\right|_{i,j}+[\mathcal{A}\phi]_{i,j}=f_{i,j}.$ (12)
@@
=${}_{i,j}=$ $-\left(2A_{1}\delta_{r}^{2}+\frac{2A_{2}}{r_{i}^{2}}\delta_{\theta}^{2}+\frac{A_{3}}{h_{r_{i-1}}}\delta_{r}+\frac{A_{4}}{r_{i}^{2}h_{\theta_{j-1}}}\delta_{\theta}\right)\phi_{i,j}+\left(\frac{A_{3}}{h_{r_{i-1}}}+A_{5}\delta_{r}+c_{1}\right)\phi_{r_{i,j}}$ (13)
$+\left(\frac{A_{4}}{r_{i}^{2}h_{\theta_{j-1}}}+\frac{A_{6}}{r_{i}^{2}}\delta_{\theta}+c_{2}\right)\phi_{\theta_{i,j}}$
where the coefficients are given as,
$A_{1}=\dfrac{2\alpha_{r}}{1+\alpha_{r}^{2}}+\dfrac{(1-\alpha_{r})^{2}}{2\alpha_{r}},$
$A_{2}=\dfrac{2\alpha_{\theta}}{1+\alpha_{\theta}^{2}}+\dfrac{(1-\alpha_{\theta})^{2}}{2\alpha_{\theta}},$
$A_{3}=\dfrac{2(1-\alpha_{r})}{(1-\alpha_{r}+\alpha_{r}^{2})}\left\{\dfrac{\alpha_{r}}{(1+\alpha_{r}^{2})}+\dfrac{(1-\alpha_{r})^{2}}{\alpha_{r}}\right\},$
$A_{4}=\dfrac{2(1-\alpha_{\theta})}{(1-\alpha_{\theta}+\alpha_{\theta}^{2})}\left\{\dfrac{\alpha_{\theta}}{(1+\alpha_{\theta}^{2})}+\dfrac{(1-\alpha_{\theta})^{2}}{\alpha_{\theta}}\right\},$
$A_{5}=\dfrac{2}{1+\alpha_{r}^{2}}$
$A_{6}=\dfrac{2}{1+\alpha_{\theta}^{2}}.$
The operators $\delta_{r}$, $\delta_{\theta}$, $\delta_{r}^{2}$ and $\delta_{\theta}^{2}$ along with Crank-Nicolson approximation yields the fully discretized form of equation 1 as
@@
$-B_{6}\left(\phi_{r_{i+1,j}}^{n+1}-\phi_{r_{i-1,j}}^{n+1}\right)-\left(B_{7}+c_{1}\right)\phi_{r_{i,j}}^{n+1}-B_{8}\left(\phi_{\theta_{i,j+1}}^{n+1}-\phi_{\theta_{i,j-1}}^{n+1}\right)-\left(B_{9}+c_{2}\right)\phi_{\theta_{i,j}}^{n+1}$
$-B_{6}\left(\phi_{r_{i+1,j}}^{n}-\phi_{r_{i-1,j}}^{n}\right)-\left(B_{7}+c_{1}\right)\phi_{r_{i,j}}^{n}-B_{8}\left(\phi_{\theta_{i,j+1}}^{n}-\phi_{\theta_{i,j-1}}^{n}\right)-\left(B_{9}+c_{2}\right)\phi_{\theta_{i,j}}^{n}$